OpenStudy (anonymous):

Write an algebraic expression which represents the volume of a cylinder whose radius is 3k and height is 2k + 1.

OpenStudy (anonymous):

\[Vc=\pi r^2h\] \[V_c=\pi (3k)^2(2k+1)\] \[V_c=9\pi(2k+1) k^2\] which is the desired algebraic expression in variable\[k\]